OpenStudy (anonymous):

using differentiation from first principles, find the solution to: y=sin(2x+1)

OpenStudy (jamesj):

You mean find the derivative, yes?, not the solution.

OpenStudy (jamesj):

and if so, the derivative in general, or at a point?

OpenStudy (anonymous):

in general

OpenStudy (jamesj):

What is your first principles definition of derivative?

OpenStudy (anonymous):

\[dy/dx=\lim_{h \rightarrow 0}[f(x+h)-f(x)]\div(h)\]

OpenStudy (jamesj):

Hence for your function this ratio is \[ \frac{\sin(2(x+h)+1) - \sin(2x+1)}{h} \] \[ = \frac{2\sin\left(\frac{2(x+h)+1 + (2x+1)}{2}\right).\cos\left(\frac{2(x+h)+1 - (2x+1)}{2}\right)}{h} \] Now you take it from here.

OpenStudy (anonymous):

please complete

OpenStudy (anonymous):

\[\int\limits_{}^{}\cos ^{4}xdx\]

OpenStudy (jamesj):

sign error. Should be \[ = \frac{2\sin([2(x+h)+1 - (2x+1)]/2).\cos([2(x+h)+1+(2x+1)]/2)}{h} \] \[ = 2\frac{\sin(h)}{h} . \cos(2x + 1 + h)\] Now take the limit of that as h --> 0.

OpenStudy (jamesj):

and remember that \[ \lim_{h \rightarrow 0} \frac{\sin h}{h} = 1 \]

OpenStudy (jamesj):

good

OpenStudy (jamesj):

;-)

OpenStudy (jamesj):

i'll answer a specific qn if i've time and interest

OpenStudy (jamesj):

where's your question?